OpenStudy (anonymous):

what is an equation of the line that passes through the point (-4,-1) and is parallel to and perpendicular to the line 2x+7y=14 ?

OpenStudy (anonymous):

I NEED HELP ASAP PLEASEE

OpenStudy (anonymous):

\[y=\frac{ 7x }{ 2 }+15\]

OpenStudy (anonymous):

you want two different equations? one for parallell and one for perpendicular? Note: parallel has same slope, and perpendicular has negative reciprocal of slope. The primary equation in slope intercept form should be y = -2/7x + 2

OpenStudy (anonymous):

So using the equation formula. (y-y1)=m(x-x1). You plug in your points from (-4,-1) and the slope you need (see above for parallel or perpendicular). Parallel. (y+1)=-2/7(x+4) : use algebra and get x and y on left of equation. Note, the +4 and +1 are due to -(-1) and -(-4). Perpendicular. Same thing in a sense, but the slope must be the negative reciprocal of the primary equation (line). (y+1)=7/2(x+4) again, use algebra and get the x and y on left side of equation.

OpenStudy (anonymous):

Therefore: Parallel: 7y+2x=-15 Perpendicular: 2y-7x=26

OpenStudy (anonymous):

need to know how to work this equation for the parallel i am not understanding fully...y=-5/3x+2...please help